# On the mechanism of black hole energy reduction in the Blandford-Znajek process

### v1

@@ 3 Steady axisymmetric MHD flows
The solution has $-u_{t}\ll\Omega_{\rm F}H_{\varphi}/(4\pi\eta)$ and $u_{\varphi}\ll H_{\varphi}/(4\pi\eta)$, as shown in the upper right panel of Figure 1, which means that the entire inflow is magnetically dominated, or nearly force-free. The matter energy $-u_{t}$ is nearly constant, that is, the matter appears to fall freely (see Section 2.1). $H_{\varphi}\approx{\rm const.}$ means that the poloidal electric current does not cross the magnetic field line. The matter does not have negative energy, $-u_{t}>0$, which is consistent with the GRMHD simulation results komissarov05 . Note that $u_{t}=u_{\mu}\xi^{\mu}$, $u_{\varphi}=u_{\mu}\chi^{\mu}$, $H_{\varphi}={}^{*}F_{\mu\nu}\xi^{\mu}\chi^{\nu}$, and $\Omega_{\rm F}=-F_{t\theta}/F_{\varphi\theta}$ are the same in the BL and KS coordinates komissarov04a .
@@
The bottom left panel of Figure 1 shows the velocity field of the inflow. Here we have assumed the poloidal magnetic field as a split-monopole type, $B^{\theta}=0$, and thus $v^{\theta}=0$. At the separation point, $v^{r}=0$ and $v^{\varphi}=\Omega_{\rm F}$. As approaching the horizon,
@@
These are the same asymptotic velocity components as a freely falling particle (Eq. 6) (see also vanputten03, ). The properties of electromagnetic fields are shown in the bottom right panel of Figure 1. $B_{\hat{r}}$ is finite, while $D_{\hat{\theta}}$ and $B_{\hat{\varphi}}$ diverge. The sign of $D_{\theta}$ is the same as that of $E_{\theta}$ at the far zone, but changes at the point of $\Omega=\Omega_{\rm F}$. One also confirms $D^{2}<B^{2}$ in the entire region. Note that the Znajek condition is approximately satisfied, $B_{\hat{\varphi}}\approx-D_{\hat{\theta}}$, at the inner magnetosonic point. [^3] In the outflow, $\hat{B}_{\varphi}\approx\hat{E}_{\theta}$ at the outer magnetosonic point (e.g. lyubarsky09, ; toma13, ). The energy density of the electromagnetic field (Eq. 12) $e<0$ inside around the inner Alfvén point (which is inside the ergosphere) and diverges.
@@ 4 Mechanism of BH energy reduction
As confirmed in the previous sections, the event horizon does not yield any condition that affects its exterior, and thus the BZ process is not like the unipolar induction process of pulsars. The matter energy $-u_{t}>0$ in the inflow, so that the BZ process is not the same as the mechanical Penrose process. As introduced in Section 1, GRMHD simulations in the KS coordinates explain that the BH reduces its rotational energy by the outward Poynting flux $S^{i}=-\alpha T^{i}_{\mu}\xi^{\mu}$ on the event horizon, but this explanation is not intuitive. Here we provide a clear, intuitive picture for the mechanism of BH energy reduction in the BZ process by using the BL coordinates.
BHs cannot have magnetic field by electric current inside themselves, unlike pulsars. The BH magnetosphere is formed by advection of the magnetic field frozen in the accretion disk (or RIAF) and disk wind, as shown in the GRMHD simulations (e.g. mckinney04, ; barkov08, ; tchekhovskoy11, ). Therefore, the past accreted plasma threaded by the magnetic field should remain just above the horizon in the BL coordinates, as considered in the membrane paradigm thorne86 and illustrated in Figure 2. This past accreted plasma keep falling towards the horizon, and we call it ‘falling membrane’. This holds waves and turbulence that have established the transverse force balance of ordered magnetic field in the magnetosphere. The magnetic field must not overtake the falling membrane, so that it does not thread the event horizon, and no Poynting flux is directly released from the horizon. In the magnetosphere with the established ordered field, the steady structure is formed between the inner and outer magnetosonic surfaces, as explained in Section 3, and the inflow which operates the BZ process is following the falling membrane.
Figure 2: Schematic picture of the generic structure of the BH magnetosphere operating the BZ process, which magnifies the region close to the horizon for illustrative purpose. The magnetic pressure of the magnetosphere makes the RIAF thin near the horizon. The past accreted plasma or ‘falling membrane’ with velocity $V<0$ is followed by the magnetically-dominated inflow (like Fig. 1) with $v^{r}<0$. The magnetic field lines thread the falling membrane but not the horizon. The outward Poynting flux $S^{r}=E_{\theta}H_{\varphi}/4\pi$ is produced at the boundary between the falling membrane and the inflow, which is associated with the inward extension of region of $e_{\rm in}<0$ as illustrated in the top right panel. The poloidal electric current $J^{r}$ just flows inwards, maintaining $H_{\varphi}$, without crossing the magnetic field lines.
@@
$D^{\theta}=D^{\theta}_{\rm mb}(r)H(-R)+D^{\theta}_{\rm in}(r)H(R),$ (30)
$H_{\varphi}=H_{\varphi}^{\rm mb}(r)H(-R)+H_{\varphi}^{\rm in}(r)H(R),$ (31)
where $H(R)$ is a Heaviside step function with $R\equiv r-r_{\rm b}$. Here and hereafter the quantities with subscript or superscript ‘mb’ represent the field components in the falling membrane, while those with ‘in’ represent the field components given by the caculation result of the steady, axisymmetric MHD inflow as shown in Figure 1. Substituting these equation into the Ampère-Maxwell equation (in Eq. 8), which gives us for $R\to 0$
$H_{\varphi}^{\rm in}-H_{\varphi}^{\rm mb}=\sqrt{\gamma}V(D^{\theta}_{\rm in}-D^{\theta}_{\rm mb})-4\pi\sqrt{\gamma}\eta^{\theta},$ (32)
where $\eta^{\theta}$ is possible surface electric current. The membrane is falling by the dominant gravitational force, so that one has (Eq. 6),
@@
Generally the relation $B_{\hat{\varphi}}\approx-D_{\hat{\theta}}$ is satisfied at $r\approx r_{\rm H}$, which can be rewritten as
@@
This relation is valid for the electromagnetic fields at both sides of the boundary, and thus rewritten by
@@
at both sides of the boundary. Therefore, Eq. (32) implies
@@
The electromagnetic energy density of the inflow (Eq. 12) at $r\approx r_{\rm H}$ is written by taking account of our assumption $B^{\theta}=0$ and Eq. (9) as (cf. komissarov09, )
$e_{\rm in}\approx\frac{1}{8\pi\alpha}\left[\alpha^{2}(B^{r}_{\rm in}B_{r}^{\rm in}+B^{\varphi}_{\rm in}B_{\varphi}^{\rm in})+\gamma_{\varphi\varphi}(\Omega_{\rm F}^{2}-\Omega_{\rm H}^{2})B^{r}_{\rm in}B_{r}^{\rm in}\right]=\frac{-\sqrt{\gamma}}{4\pi}\Omega_{\rm F}D^{\theta}_{\rm in}B_{\rm in}^{r}<0,$ (37)
@@
Therefore, the Poynting flux is produced at the boundary between the falling membrane and inflow, not at the event horizon, and its production is associated with the inward extension of region with $e_{\rm in}<0$, as illustrated in the top right panel of Figure 2. The creation of $e_{\rm in}<0$ is equivalent to reduction of the rotational energy of the Kerr spacetime. We note that $e_{\rm in}$ diverges but $V$ converges to zero. The sign of the electromagnetic energy density of the falling membrane can be either positive or negative, which is not relevant for the mechanism.
@@ 5 Discussion
In this paper, we have treated the falling membrane that should exist above the event horizon in the BL coordinates and clarified the generic existence of non-steady front of the inflow where the outward Poynting flux and negative electromagnetic energy are created (see Eq. 38). Most of the previous studies except the membrane paradigm thorne86 assume that the magnetic field lines thread the event horizon even in the BL coordinates (e.g. blandford77, ; takahashi90, ; takahashi08, ; beskin00, ; huang20, ). The situation without the failling membrane is simple for searching the MHD flow solutions, but unrealistic and not suitable for understanding the mechanism of BH energy reduction.
@@ 5.1 Previous studies on the BH energy reduction
The membrane paradigm considers that the magnetic field lines are threading the membrane of past accreted matter similarly to that illustrated in Figure 2 thorne86 . Since the matter falls with velocity $v^{r}\approx-\alpha/\sqrt{\gamma_{rr}}\to 0$ for $r\to r_{\rm H}$, the membrane paradigm assumes that the membrane is at a fixed radius, i.e., $V=0$, and that the conductive surface electric current $\eta^{\theta}=\alpha D^{\theta}/4\pi\sqrt{\gamma_{rr}}$ produces $H_{\varphi}^{\rm in}$. In reality, however, the membrane keeps falling with convergently small velocity, and $\eta^{\theta}\approx 0$ as shown in Section 4. Inside the membrane the electromagnetic field is arbitrary, depending on the past accretion history (see also punsly89, ). $H_{\varphi}$ and $\Omega_{\rm F}$ are determined by the frame dragging effect between the inner and outer magnetosonic surfaces. The membrane paradigm is misleading with respect to the origin of the Poynting flux and the structure of poloidal electric current.
Invariant properties in the BL and KS coordinates

### v2

@@ 3 Steady axisymmetric MHD flows
The solution has $-u_{t}\ll-H_{\varphi}\Omega_{\rm F}/4\pi|\eta|$ and $u_{\varphi}\ll-H_{\varphi}/4\pi|\eta|$, as shown in the upper right panel of Figure 1, which means that the entire inflow is magnetically dominated. The matter energy $-u_{t}$ is nearly constant, that is, the matter appears to fall freely (see Section 2.1). $H_{\varphi}\approx{\rm const.}$ means that the poloidal electric current does not cross the magnetic field line (see Section 2.3). The matter does not have negative energy, $-u_{t}>0$, which is consistent with the GRMHD simulation results komissarov05 . Note that $u_{t}=u_{\mu}\xi^{\mu}$, $u_{\varphi}=u_{\mu}\chi^{\mu}$, $H_{\varphi}={}^{*}F_{\mu\nu}\xi^{\mu}\chi^{\nu}$, and $\Omega_{\rm F}=-F_{t\theta}/F_{\varphi\theta}$ are the same in the BL and KS coordinates komissarov04a .
@@
The bottom left panel of Figure 1 shows the velocity field of the inflow. At the separation point, $v^{r}=0$ and $v^{\varphi}=\Omega_{\rm F}$. As approaching the horizon,
@@
These are the same asymptotic velocity components as a freely falling particle (Eq. 6) (see also vanputten03, ). The properties of electromagnetic fields are shown in the bottom right panel of Figure 1. $B_{\hat{r}}$ is finite, while $D_{\hat{\theta}}$ and $B_{\hat{\varphi}}$ diverge. The sign of $D_{\theta}$ is the same as that of $E_{\theta}$ at the far zone, but changes at the point of $\Omega=\Omega_{\rm F}$. One also confirms $D^{2}<B^{2}$ in the entire region. Note that the Znajek condition is approximately satisfied, $B_{\hat{\varphi}}\approx-D_{\hat{\theta}}$, at the inner magnetosonic point. [^3] In the outflow, $B_{\hat{\varphi}}\approx E_{\hat{\theta}}$ at the outer magnetosonic point (e.g. lyubarsky09, ; toma13, ). The energy density of the electromagnetic field (Eq. 12) $e<0$ inside around the inner Alfvén point and diverges. We note that the inner Alfvén point is close to the inner light surface, $f=0$, which is generically inside the ergosphere komissarov04a , and $e$ can be negative only inside the ergosphere komissarov09 .
Actually our solution of $B_{\hat{r}}$ deviates from the assumed monopole shape, but we confirm that the deviation is $<12\%$ ($<18\%$ in the solution for $\theta=\pi/6$). Adjusting the assumed $\mathbf{B}_{\rm p}$ is required for obtaining exact solutions, but we consider that our solution is sufficient for understanding the detailed flow structure near the horizon and providing with the picture for the BH energy reduction mechanism (Section 4).
Figure 2: Characteristic radii of our MHD inflow solutions as a function of $\mathcal{E}$. For the top left panel, the parameter sets $(a,\Omega_{\rm F}/\Omega_{\rm H},\mathcal{E}-\mathcal{L}\Omega_{\rm F})$ are $(0.8,0.5,0.55)$, which include the case of Figure 1. We plot the radii of separation point (thick solid line), inner Alfvén point (square), inner magnetosonic point (black dot), point of $e=0$ (thin solid line), that of $|v^{r}|/(\alpha/\sqrt{\gamma_{rr}})=0.99$ (dashed line), and that of $D_{\hat{\theta}}/|B_{\hat{\varphi}}|=0.99$ (dot-dashed line) for the case of each $\mathcal{E}$. The parameter sets for the top right, bottom left, and bottm right panels are $(a,\Omega_{\rm F}/\Omega_{\rm H},\mathcal{E}-\mathcal{L}\Omega_{\rm F})=(0.5,0.5,0.55)$, $(0.8,0.4,0.55)$, and $(0.8,0.5,0.45)$, respectively.
We also obtain solutions with other parameter value, $a=0.5$, $\Omega_{\rm F}/\Omega_{\rm H}=0.4$, or $\mathcal{E}-\mathcal{L}\Omega_{\rm F}=0.45$ to confirm that they exhibit similar structures of inflows (Figure 2). $\Omega_{\rm F}/\Omega_{\rm H}=0.4$ is within the range, $\Omega_{\rm F}/\Omega_{\rm H}\sim 0.3-0.5$, which the GRMHD simulations indicate tchekhovskoy10 ; penna13 . For a given $\Omega_{\rm F}$, one has the maximum value of $-f$ from Eq. (23), and then the existence of the separation point ($u_{\rm p}^{2}=0$ in Eq. 22) requires $\mathcal{E}-\mathcal{L}\Omega_{\rm F}<0.58$ for $(a,\Omega_{\rm F}/\Omega_{\rm H})=(0.8,0.5)$. The cases of $(a,\Omega_{\rm F}/\Omega_{\rm H})=(0.8,0.4)$ and $(0.5,0.5)$ have similar constraints, $\mathcal{E}-\mathcal{L}\Omega_{\rm F}<0.64$ and $<0.72$, respectively. Figure 2 plots the radii of separation point (thick solid line), inner Alfvén point (square), inner magnetosonic point (black dot), point of $e=0$ (thin solid line), that of $|v^{r}|/(\alpha/\sqrt{\gamma_{rr}})=0.99$ (dashed line), and that of $D_{\hat{\theta}}/|B_{\hat{\varphi}}|=0.99$ (dot-dashed line) for the case of each $\mathcal{E}$. $|\mathcal{E}|\gg 1$ is required for magnetically-dominated flows. As can be seen in Figure 2, $e<0$ inside around the inner Alfvén point, and $B_{\hat{\varphi}}\approx-D_{\hat{\theta}}$ around the inner magnetosonic point. The ergosphere is inside the radius of $r_{\rm es}-r_{\rm H}\approx 0.35$ for $a=0.8$ and $\approx 0.1$ for $a=0.5$, so that all the characteristic radii except the separation point are in the ergosphere. We also confirm the same properties for the solutions with $\theta=\pi/6$ (see Appendix B).
@@ 4 Mechanism of BH energy reduction
As confirmed in Section 3, the matter energy $-u_{t}>0$ in the inflow, so that the BZ process is not the same as the mechanical Penrose process. Instead, the electromagnetic energy density (in the BL coordinates) $e<0$ inside around the inner Alfvén surface. Focusing on these properties, some theorists tried to find a picture how the BH reduces its rotational energy with infall of negative electromagnetic energy, but all of the proposed pictures were ambiguous komissarov09 ; takahashi90 ; koide14 (see Section 5.1.3). Below we provide a clear picture for the BH energy reduction by the BZ process in the BL coordinates.
BHs cannot have magnetic field by electric current inside themselves, unlike pulsars. The BH magnetosphere is formed by advection of the magnetic field frozen in the accretion disk (or RIAF) and disk wind, as shown in the GRMHD simulations (e.g. mckinney04, ; barkov08, ; tchekhovskoy11, ). Therefore, the past accreted plasma threaded by the magnetic field should remain just above the horizon in the BL coordinates, as considered in the membrane paradigm thorne86 and illustrated in Figure 3. This past accreted plasma keeps falling towards the horizon, and we call it ‘falling membrane’. This holds waves and turbulence that have established the transverse force balance of ordered magnetic field in the magnetosphere. The magnetic field must not overtake the falling membrane under the MHD condition, so that it does not thread the event horizon, and no Poynting flux is directly released from the horizon. In the magnetosphere with the established ordered field, the steady structure is formed between the inner and outer magnetosonic surfaces, as explained in Section 3, and the inflow $(v^{r}<0)$ which carries the electromagnetic field $E_{\theta}$ and $H_{\varphi}$ (forming the outward Poynting flux $S^{r}=E_{\theta}H_{\varphi}/4\pi\sqrt{\gamma}$) is following the falling membrane. The inflow inside the inner magnetosonic surface keeps extending towards the event horizon, and in this sense, the front of the inflow is not steady.
Figure 3: Schematic picture of the generic structure of the BH magnetosphere operating the BZ process, which magnifies the region close to the horizon for illustrative purpose. The magnetic pressure of the magnetosphere makes the RIAF thin near the horizon. The past accreted plasma or ‘falling membrane’ with velocity $V<0$ is followed by the magnetically-dominated inflow (like Fig. 1) with $v^{r}<0$. The magnetic field lines thread the falling membrane but not the horizon. The outward Poynting flux $S^{r}=E_{\theta}H_{\varphi}/4\pi\sqrt{\gamma}$ is produced at the boundary between the falling membrane and the inflow, which is associated with the inward extension of region of $e_{\rm in}<0$ as illustrated in the top right panel. The poloidal electric current $J^{r}$ just flows inwards, maintaining $H_{\varphi}$, without crossing the magnetic field lines.
@@
$D^{\theta}=D^{\theta}_{\rm mb}H(-R)+D^{\theta}_{\rm in}H(R),$ (27)
$H_{\varphi}=H_{\varphi}^{\rm mb}H(-R)+H_{\varphi}^{\rm in}H(R),$ (28)
where $H(R)$ is a Heaviside step function with $R\equiv r-r_{\rm b}$. Here and hereafter the quantities with subscript or superscript ‘mb’ represent the field components in the falling membrane, while those with ‘in’ represent the field components given by the calculation result of the steady, axisymmetric inflow as shown in Figure 1. Substituting these equation into the Ampère-Maxwell equation (in Eq. 8) leads to
$-D^{\theta}_{\rm mb}V\delta(R)+D^{\theta}_{\rm in}V\delta(R)-\frac{1}{\sqrt{\gamma}}(\partial_{r}H_{\varphi}^{\rm mb})H(-R)+\frac{1}{\sqrt{\gamma}}H_{\varphi}^{\rm mb}\delta(R)$
$+\frac{1}{\sqrt{\gamma}}(\partial_{r}H_{\varphi}^{\rm in})H(R)-\frac{1}{\sqrt{\gamma}}H_{\varphi}^{\rm in}\delta(R)=4\pi\eta^{\theta}\delta(R),$ (29)
where $\eta^{\theta}$ is possible surface electric current. This gives us for $R\to 0$
$H_{\varphi}^{\rm in}-H_{\varphi}^{\rm mb}=\sqrt{\gamma}V(D^{\theta}_{\rm in}-D^{\theta}_{\rm mb})-4\pi\sqrt{\gamma}\eta^{\theta}.$ (30)
The membrane is falling by the dominant gravitational force, so that one has (Eq. 6),
@@
Generally the relation $B_{\hat{\varphi}}\approx-D_{\hat{\theta}}$ is satisfied at $r\approx r_{\rm H}$ (Eq. 14), which can be rewritten as
@@
Thus this relation is valid for the electromagnetic fields in the falling membrane. We have seen that this relation is also valid inside around the inner magnetosonic surface in the inflow (Section 3). Therefore, by using Eq. (31) valid for both sides of the boundary, we obtain
@@
at both sides of the boundary. This implies in Eq. (30)
@@
The electromagnetic energy density of the inflow (Eq. 12) at $r\approx r_{\rm H}$ is written by taking account of our assumption $B^{\theta}=0$ and Eq. (9) (cf. komissarov09, ) as well as the condition inside around the inner magnetosonic surface, $B^{r}_{\rm in}B_{r}^{\rm in}\ll B^{\varphi}_{\rm in}B_{\varphi}^{\rm in}\approx D^{\theta}_{\rm in}D_{\theta}^{\rm in}$,
$e_{\rm in}$ $=\frac{1}{8\pi\alpha}\left[\alpha^{2}(B^{r}_{\rm in}B_{r}^{\rm in}+B^{\varphi}_{\rm in}B_{\varphi}^{\rm in})+\gamma_{\varphi\varphi}(\Omega_{\rm F}^{2}-\Omega^{2})B^{r}_{\rm in}B_{r}^{\rm in}\right]$ (35)
$\approx\frac{\gamma_{rr}\gamma_{\varphi\varphi}}{4\pi\alpha}\Omega_{\rm F}(\Omega_{\rm F}-\Omega_{\rm H})(B^{r}_{\rm in})^{2}=\frac{-\sqrt{\gamma}}{4\pi}\Omega_{\rm F}D^{\theta}_{\rm in}B_{\rm in}^{r}<0,$
@@
Therefore, the Poynting flux is produced at the boundary between the falling membrane and the inflow, not at the event horizon, and its production is associated with the inward extension of region with $e_{\rm in}<0$, as illustrated in the top right panel of Figure 3. The creation of $e_{\rm in}<0$ is equivalent to reduction of the rotational energy of the Kerr spacetime. We note that $e_{\rm in}$ diverges but $V$ converges to zero. The sign of the electromagnetic energy density of the falling membrane can be either positive or negative, which is not relevant for the mechanism of BH energy reduction.
In our picture, the ergosphere plays the role of making $e_{\rm in}<0$. Let us recall that the BZ process is the steady production of the outward Poynting flux along the magnetic field lines threading the horizon. This steady process is established between the inner and outer magnetosonic surfaces. Thus the front of the inflow is already inside the inner magnetosonic surface, where $e_{\rm in}<0$, since the inner magnetosonic surface is well inside the ergosphere. Consequently, the spacetime reduces its rotational energy by the inward movement of the inflow front.
@@ 5 Discussion
In this paper, we have treated the membrane of past accreted matter that should exist above the event horizon and falls with velocity $V\approx-\alpha/\sqrt{\gamma_{rr}}$ in the BL coordinates. We have clarified the generic existence of unsteady front of the magnetically-dominated inflow where the outward Poynting flux $S_{\rm in}^{r}=E_{\theta}^{\rm in}H_{\varphi}^{\rm in}/4\pi\sqrt{\gamma}$ and negative electromagnetic energy $e_{\rm in}$ are created (see Eq. 36). The creation of $e_{\rm in}<0$ is equivalent to the rotational energy reduction of the Kerr spacetime. Most of the previous studies except the membrane paradigm thorne86 assume that the magnetic field lines thread the event horizon even in the BL coordinates (e.g. blandford77, ; takahashi90, ; takahashi08, ; beskin00, ; huang20, ). The situation without the falling membrane is simple for searching the MHD flow solutions, but unrealistic and not suitable for understanding the mechanism of BH energy reduction.
@@ 5.1.1 The membrane paradigm
The membrane paradigm considers that the magnetic field lines are threading the membrane of past accreted matter similarly to that illustrated in Figure 3 thorne86 . Since the matter falls with velocity $V\approx-\alpha/\sqrt{\gamma_{rr}}\to 0$ for $r\to r_{\rm H}$, the membrane paradigm assumes that the membrane is at a fixed radius, i.e., $V=0$, and that the origin of $H_{\varphi}^{\rm in}$ is conductive surface electric current,
$\sqrt{\gamma_{rr}}\eta^{\theta}=\frac{1}{4\pi}\alpha D^{\theta}.$ (37)
Indeed, Eq. (37) is derived from Eqs. (30) and (32) by setting $V=0$ and $H_{\varphi}^{\rm mb}=0$. In the paradigm, Eq. (37) is interpreted as Ohm’s law, and the conductive surface current is considered to close the poloidal electric current circuit and to reduce the BH energy by $\mathbf{J}\times\mathbf{B}$ force.
In reality, however, the membrane keeps falling with convergently small velocity $V$, and $\eta^{\theta}\approx 0$ as shown in Section 4. $H^{\rm in}_{\varphi}$ is not formed by the cross-field current $\eta^{\theta}$, but determined by the conditions at the inner and outer magnetosonic surfaces (or the frame-dragging effect between the surfaces), as explained in Section 3. Inside the membrane the electromagnetic field is arbitrary, depending on the past accretion history (see also punsly89, ). The membrane paradigm is misleading with respect to the origin of the Poynting flux, the BH energy reduction mechanism, and the structure of poloidal electric current.
5.1.2 Different views in the BL and KS coordinates